OpenStudy (anonymous):

Which is the equation for the line perpendicular to y=-5/3x+11 1/3 and containing P(-2,3). How do i solve this one?? :(

OpenStudy (anonymous):

im sorry i cant figure this one out

OpenStudy (anonymous):

it's okay. i wish i could :(

OpenStudy (anonymous):

here try this website www.library.pima.gov and click on the home work help and it will ask u what subject you need help with and ur grade...tell them tht and then they will ask you to give them a code you can use mine....its 21152028698797 and this should help you out with ur homework and studies. Hope it helps.

OpenStudy (anonymous):

thanks so much. i'll try it.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Did tht help you???

OpenStudy (anonymous):

yep. It did. Thanks so much! :)

OpenStudy (anonymous):

you are welcome and you can use tht when ever u want...i hardly use it

OpenStudy (anonymous):

Cool and thanks again. I think i'll be needing it alot until skool's out.

OpenStudy (anonymous):

Slope= 3/5 (opp reciprocal) y=mx+b 3=3/5(-2)+b 3=-6/5+b 3+6/5=b 4.2=b y=1/5x+4.2

OpenStudy (anonymous):

thank you x3pufflix3

OpenStudy (anonymous):

Anytime